# Smoothing of Headland Path Edges and Headland-to-Mainfield Lane Transitions Based on a Spatial Domain Transformation and Linear Programming

### v3

@@ Abstract
Within the context of in-field path planning and under the assumption of nonholonomic vehicle models this paper addresses two tasks: smoothing of headland path edges and smoothing of headland-to-mainfield lane transitions. Both tasks are solved by a two-step hierarchical algorithm. The first step differs for the two tasks generating either a piecewise-affine or a Dubins reference path. The second step leverages a transformation of vehicle dynamics from the time domain into the spatial domain and linear programming. Benefits such as a hyperparameter-free objective function and spatial constraints useful for area coverage gaps avoidance and precision path planning are discussed. The method, which is a deterministic optimisation-based method, is evaluated on 5 real-world fields solving 19 instances of the first task and 84 instances of the second task.
@@
In-field path planning; Nonholonomic vehicle dynamics; Linear Programming; Spatial Transformation.
@@ 1 Introduction
MAIN NOMENCLATURE
@@
$e_{\psi}$ Path-aligned relative heading, (deg).
@@
$\delta$ Steering angle of vehicle, (deg).
$\psi$ Global absolute heading, (deg).
@@
Abbreviations
LP Linear Programming.
PWA Piecewise-Affine.
UTM Universal Transverse Mercator coordinate system.
Logistics play a large role in agriculture. For arable farming and the cultivation of crops it can be differentiated between in-field and inter-field logistics (Plessen (2019a)). For the former, efficient path planning is of interest subject to three main considerations: (i) minimisation of tractor traces within the field in order to maximise growable field area, (ii) minimisation of area coverage pathlength in order to minimise fuel consumption and time spent in the field, and (iii) avoiding area coverage gaps (unsprayed areas) such that the entire field is covered, for example during a field-run spraying herbicide or pesticide applications, while simultaneously accounting for a given operating width of the machinery operating in the field and accounting for its typically nonholonomic vehicle dynamics, state and actuator constraints.
Within this context the motivation and contribution of this paper is to present a method for smoothing of headland path edges and headland-to-mainfield lane transitions based on a spatial domain transformation and linear programming. As will be motivated below, the spatial domain transformation permits to formulate spatial constraints useful for precision path planning in agriculture. A linear programming framework with a hyperparameter-free objective function can incorporate those constraints together with linearised and discretised nonholonomic vehicle dynamics in a disciplined manner to produce deterministic precision paths and fast solve times.
@@
Thus, given an optimised sequence of mainfield lane traversals for field coverage such as produced in Plessen (2019b) and Plessen (2018), this paper provides a method for path smoothing. See also Höffmann et al. (2024), Mier et al. (2023), Khan et al. (2016) for more high-level area coverage techniques. This paper differs from Plessen & Bemporad (2017), where an online closed-loop tracking controller based on model predictive control was presented, in that here the focus is on offline generation of precision paths, which might then afterwards be tracked online by the approach from aforementioned paper or an alternative closed-loop tracking controller.
@@
A widespread approach to reference path and in particular headland turn trajectory generation is to build methods around specific path planning patterns. A literature review is provided. In Trendafilov & Tihanov (2022) 5 types of T-turns are considered. In Wang & Noguchi (2018) adaptive turning based on a switchback-pattern and dynamic circleback-pattern is discussed. In Sabelhaus et al. (2013) 9 different symmetric lane-to-lane transitions are analysed, including Omega-turn, U-turn, gap-turn and fishtail-turn. Similarly, Backman et al. (2015) account for maximum steering rate and the maximum acceleration of the vehicle. In Bulgakov et al. (2019) for a trailed asymmetric swath reaper–tractor aggregate a pear-shaped loop-turn and a U-turn are discussed. In Boryga et al. (2020) 4 types of patterns are discussed, whereby each is composed of two transition curves. In Peng et al. (2023) an optimisation‐based method for turning in constrained headland spaces such as orchards is presented, using 4 types of turns (U-turn, Omega-turn, switchback-turn and circleback-turn) for generation of a first reference path. Gao et al. (2023) use Bezier curves for a headland-turning navigation system for a safflower picking robot. Finally, see Ha et al. (2018) for an example of a hardware modification to facilitate headland turning.
@@
Once a smooth path plan is computed, for autonomous robots a plethora of techniques for online tracking can be applied, see e.g. Wang et al. (2023), Ettefagh et al. (2018), He et al. (2023), Wang et al. (2024), Sun et al. (2024), Liu et al. (2024), Eski & Kuş (2019), Jing et al. (2021), Yin et al. (2020). For a recent overview of control algorithms of autonomous all-terrain vehicles in agriculture see Etezadi & Eshkabilov (2024).
@@
The research gap and motivation for this paper is discussed. First, the focus of aforementioned pattern-based references is on lane-to-lane transitions. Second, an issue for pattern-based path planning, in particular based on straights, arcs and clothoid segments, is that resulting paths are typically operating at actuation limits such as maximum curvature and maximum steering rate. For general headland path smoothing (in contrast to aforementioned lane-to-lane transitions) this is not sufficient where other constraints such as area coverage gaps must also be accounted for. Third, the proposal of a spatial domain transformation in order to be able to formulate optimisation problems with spatial constraints, which is desirable for precision agriculture applications, is missing in the literature.
@@
The focus of this paper is on path planning instead of path tracking. The allure of a numerical optimisation approach is that it allows incorporation of a variety of constraints in a disciplined manner and can handle the multivariate nature of the task. The allure of transforming vehicle dynamics from the time into the spatial domain is that it allows to formulate spatial precision constraints. An early paper discussing benefits of transformations of dynamics from the time into the spatial domain is Pfeiffer & Johanni (1987). To the author’s knowledge this is the first paper that proposes a spatial transformation of system dynamics to calculate precision path plans in agriculture.
@@
The remaining paper is organised as follows: problem formulation, modeling and proposed solution, numerical results and the conclusion are described in Sections 2-6.
2 Problem Formulation
@@
Given headland-to-mainfield lane transitions, and vice versa mainfield lane-to-headland transitions, as part of a field covering path, correct those if necessary such that smoothed transitions are generated that (i) are compatible with nonholonomic vehicle dynamics and (ii) avoid any area coverage gaps.
@@
3 Proposed Solution
The general approach to address both Problems 1 and 2 is a hierarchical two-step algorithm. The first step differs for the two problems and has as objective the generation of a reference path. The second step uses that reference path and leverages linear programming and a transformation of vehicle dynamics from the time into the spatial domain.
@@ 3 Proposed Solution
Fig. 3: The kinematic bicycle model including the representation of the path-aligned coordinate system.
@@ 3.1 Spatial Domain Transformation
To illustrate the spatial domain transformation and for simplicity, the method is illustrated for a kinematic bicycle model as shown in Fig. 3. It is referred to Plessen et al. (2017a) for handling of more complex vehicle models.
This section summarises the notation of Plessen et al. (2017b). Consider a global and a path-aligned coordinate frame within the $(x,y)$- and $(s,e_{y})$-plane, respectively, with path centerline coordinate $s\geq 0$ and lateral deviation $e_{y}\in\mathbb{R}$. Coordinates can be transformed by projecting point masses (vehicle’s center of gravity) to piecewise-affine path segments. As a detail, it can be differentiated between the path centerline coordinate $s\geq 0$, and actual path length coordinate $\eta\geq 0$, indicating the distance traveled by the vehicle. Note that $\eta\neq s$, unless the vehicle is traveling perfectly along the path centerline. A trajetory can then be defined as $z(s)=\begin{bmatrix}e_{\psi}(s),~{}e_{y}(s)\end{bmatrix}$, with heading deviation $e_{\psi}(s)\in[0,2\pi]$ at coordinate $s$. The equivalent trajectory in the $(x,y)$-plane is defined by $\mathcal{X}(s)=\begin{bmatrix}x(s),~{}y(s),~{}\psi(s)\end{bmatrix}$. The classic nonlinear kinematic bicycle model Rajamani (2011) is
@@
assuming the center of gravity to be located at the rear axle and $l$ denoting the wheelbase. The front-axle steering angle is $\delta$ and vehicle velocity is $v$. Let us denote time and spatial derivatives by $\dot{x}=\frac{dx}{dt}$ and $x^{\prime}=\frac{dx}{ds}$, respectively. The spatial equivalent of (1) is derived as follows. By $\dot{e}_{\psi}=\dot{\psi}-\dot{\psi}_{s}$ where $\psi_{s}\in[0,2\pi]$ the path heading, $\dot{e}_{y}=v\sin(e_{\psi})$, $\dot{s}=\frac{\rho_{s}v\cos(e_{\psi})}{\rho_{s}-e_{y}}$ where $\rho_{s}$ denotes the radius of curvature, $e_{\psi}^{\prime}=\frac{\dot{e}_{\psi}}{\dot{s}}$ and $e_{y}^{\prime}=\frac{\dot{e}_{y}}{\dot{s}}$, we obtain
@@
Importantly, (2) is entirely independent of vehicle speed $v$. This is characteristic for kinematic models, but not the case for dynamic vehicle models (Plessen et al. (2017a)). Furthermore, note that (2) has one state dimension less than (1). To summarise, we define spatially-dependent state and control vectors as $z=\begin{bmatrix}e_{\psi},~{}e_{y}\end{bmatrix}$ and $u=\begin{bmatrix}\delta\end{bmatrix}$, respectively. We abbreviate (2) by $z^{\prime}=f(z,u)$. Let a discretisation grid along the path centerline be defined by $\{s_{j}\}_{j=0}^{N}=\{s_{0},s_{1},\dots,s_{N}\}$. For the remainder of this paper, index $j$ refers to the spatial discretisation grid. The discretisation grid can be uniformly spaced. Additional grid points can be added. For a given set of references $\{e_{\psi,j}^{\text{ref}}\}_{j=0}^{N}$, $\{e_{y,j}^{\text{ref}}\}_{j=0}^{N}$ and $\{u_{j}^{\text{ref}}\}_{j=0}^{N-1}$ and initial state, the linearised and discretised (zero-order hold) state dynamics of (2) can be computed for $\{e_{\psi,j}\}_{j=0}^{N}$ and $\{e_{y,j}\}_{j=0}^{N}$ as a function of control inputs $\{\delta_{j}\}_{j=0}^{N-1}$.
@@
whereby the bounds, $\dot{\delta}^{\text{min}}$ and $\dot{\delta}^{\text{max}}$, in general, are nonlinear functions of the vehicle’s operating point and time-varying parameters. By applying the spatial coordinate transformation $\delta^{\prime}=\frac{\dot{\delta}}{\dot{s}}$, a discretisation, a change of variables, and assuming the bounds to remain constant over the planning horizon, we obtain
@@
whereby we abbreviated $D_{s,j}=s_{j+1}-s_{j}$. Thus, linear rate constraints in (3) are rendered not only nonlinear, but also state-dependent and also velocity-dependent. This has two implications. First, to formulate linearly constrained optimisation problems, in general, we require the linearisation of (4). Dependent on the quality of underlying reference trajectories, this may incur significant distortions. Second, while the discrete form of (3) can always be guaranteed to be feasible (assuming a feasible initialisation), for (4) this is not the case anymore. Let us consider two dregrees of simplification of (4). First, we assume given reference trajectories that permit small errors $e_{\psi,j}$ and $e_{y,j}$, and consequently approximate $\frac{(\rho_{s,j}-e_{y,j})}{\rho_{s,j}\cos(e_{\psi,j})}\approx 1$, thereby rendering (4) state-independent, but maintaining velocity-dependent bounds. Thus, steering rate constraints still depend on $v_{j}$. Second, we additionally eliminate this velocity-dependency and formulate
@@
whereby, $v_{\text{ref}}>0$ is introduced as a reference velocity parameter choice. Note that approximation (5) is the exact equivalent of (4) for the zero-reference error setting $e_{y,j}=0$, $e_{\psi,j}=0$, $v_{j}=v_{\text{ref}}$. Formulation (5) bears the advantage of eliminating state dependency and is therefore our preferred form for spatial rate constraints. In practice, we employ bounds that are constant over the spatial planning horizon, but in general time-varying in any closed-loop receding horizon control-setting and dependent on the vehicle’s operating point.
To summarise this section, the transformation of time-dependent control rate constraints (3) to the path-aligned coordinate frame is not trivial and to be considered as the main disadvantage of a spatial-based system representation. We opted for the simple state-independent form (5) for linearly constrained optimisation problems, and discussed the role of $v_{\text{ref}}>0$ as a transformation parameter.
Finally, we draw an important analogy to clothoid and continuous curvature path planning methods to further justify above approach for the handling of rate constraints in the spatial domain. Continuous curvature and clothoid based path planning methods such as in Sabelhaus et al. (2013) typically assume bounds on ’curvature’ and ’sharpness’ (the change in curvature) resulting in piecewise-affine curvature profiles along the planned path. Note that this implicitly also assumes an approximation of steering rate constraints.

### v4

@@ Abstract
Within the context of in-field path planning, and under the assumption of nonholonomic vehicle models, this paper addresses two tasks: smoothing of headland path edges and smoothing of headland-to-mainfield lane transitions. Both tasks are solved by a two-step hierarchical algorithm. The first step differs for the two tasks generating either a piecewise-affine or a Dubins reference path. The second step leverages a transformation of vehicle dynamics from the time domain into the spatial domain and linear programming. Benefits, such as a hyperparameter-free objective function and spatial constraints useful for area coverage gaps avoidance and precision path planning, are discussed. The method, which is a deterministic optimisation-based method, is evaluated on 5 real-world fields solving 19 instances of the first task and 84 instances of the second task.
@@
In-field path planning; Nonholonomic vehicle dynamics; Dubins path; Trajectory generation; Area coverage gap avoidance.
@@ 1 Introduction
NOMENCLATURE TABLE
Abbreviations
LP Linear Programming.
PWA Piecewise-Affine.
UTM Universal Transverse Mercator coordinate system.
@@
$e_{\psi}$ Path-aligned relative heading, ().
@@
Greek Symbols
$\delta$ Steering angle of vehicle, ().
$\psi$ Global absolute heading, ().
@@
Logistics play a large role in agriculture. For arable farming and the cultivation of crops it can be differentiated between in-field and inter-field logistics (Plessen (2019a)). For the former, efficient path planning is of interest subject to three main considerations: (i) minimisation of tractor traces within the field in order to maximise growable field area, (ii) minimisation of area coverage pathlength in order to minimise fuel consumption and time spent in the field, and (iii) avoiding area coverage gaps so that the entire field is covered, for example during a field-run spraying herbicide or pesticide applications, while simultaneously accounting for a given operating width of the machinery and accounting for its typically nonholonomic vehicle dynamics, state and actuator constraints.
Within this context the motivation and contribution of this paper is to present a method for smoothing of headland path edges and headland-to-mainfield lane transitions based on a spatial domain transformation and linear programming. As will be motivated below, the spatial domain transformation permits the formulation of spatial constraints useful for precision path planning in agriculture. A linear programming framework with a hyperparameter-free objective function can incorporate those constraints together with linearised and discretised nonholonomic vehicle dynamics in a disciplined manner to produce deterministic precision paths and fast solution times.
@@
Thus, given an optimised sequence of mainfield lane traversals for field coverage, such as produced in Plessen (2019b) and Plessen (2018), this paper provides a method for path smoothing. See also Höffmann et al. (2024), Mier et al. (2023), Khan et al. (2016) for more high-level area coverage techniques. This paper differs from Plessen & Bemporad (2017), where an online closed-loop tracking controller based on model predictive control was presented. In this work, the focus is on offline generation of precision paths, which might then afterwards be tracked online by the approach from aforementioned paper or an alternative closed-loop tracking controller.
@@
A widespread approach to reference path generation, and in particular headland turn trajectory generation, is to build methods around specific path planning patterns. In Trendafilov & Tihanov (2022), 5 types of T-turns are considered. In Wang & Noguchi (2018) adaptive turning based on a switchback-pattern and dynamic circleback-pattern is discussed. In Sabelhaus et al. (2013), 9 different symmetric lane-to-lane transitions are analysed, including Omega-turn, U-turn, gap-turn and fishtail-turn. Similarly, Backman et al. (2015) accounted for maximum steering rate and the maximum acceleration of the vehicle. In Bulgakov et al. (2019), for a trailed asymmetric swath reaper–tractor aggregate a pear-shaped loop-turn and a U-turn are discussed. In Boryga et al. (2020), 4 types of patterns are discussed, whereby each is composed of two transition curves. In Peng et al. (2023), an optimisation‐based method for turning in constrained headland spaces, such as orchards, is presented, using 4 types of turns (U-turn, Omega-turn, switchback-turn and circleback-turn) for generation of a first reference path. Gao et al. (2023) used Bezier curves for a headland-turning navigation system for a safflower picking robot. Finally, see Ha et al. (2018) for an example of a hardware modification to facilitate headland turning.
@@
Once a smooth path plan is computed, for autonomous robots, a plethora of techniques for online tracking can be applied, see e.g. Wang et al. (2023), Ettefagh et al. (2018), He et al. (2023), Wang et al. (2024), Sun et al. (2024), Liu et al. (2024), Eski & Kuş (2019), Jing et al. (2021), Yin et al. (2020). For a recent overview of control algorithms of autonomous all-terrain vehicles in agriculture see Etezadi & Eshkabilov (2024).
@@
The research gap and motivation for this paper is as follows. First, the focus of aforementioned pattern-based references is on lane-to-lane transitions. Second, an issue for pattern-based path planning, in particular based on straights, arcs and clothoid segments, is that resulting paths are typically operating at actuation limits, such as maximum curvature and maximum steering rate. For general headland path smoothing (in contrast to aforementioned lane-to-lane transitions), this is not sufficient where other constraints, such as area coverage gaps, must also be accounted for. Third, the proposal of a spatial domain transformation in order to be able to formulate optimisation problems with spatial constraints, which is desirable for precision agriculture applications, is missing in the literature.
@@
The focus of this paper is on path planning instead of path tracking. The allure of a numerical optimisation approach is that it allows incorporation of a variety of constraints in a disciplined manner and can handle the multivariate nature of the task. The advantage of transforming vehicle dynamics from the time into the spatial domain is that it allows to formulate spatial precision constraints. An early paper discussing benefits of transformations of dynamics from the time into the spatial domain is Pfeiffer & Johanni (1987). To the author’s knowledge this is the first paper that proposes a spatial transformation of system dynamics to calculate precision path plans in agriculture.
@@
The rest of the paper is organised as follows: problem formulation, modeling and proposed solution, numerical results and the conclusion are described in Sections 2-7.
2 Problem Formulation and Solution
@@
Given headland-to-mainfield lane transitions, and vice versa mainfield lane-to-headland transitions, as part of a field covering path, correct these if necessary such that smoothed transitions are generated that (i) are compatible with nonholonomic vehicle dynamics and (ii) avoid any area coverage gaps.
@@
The general approach to address both Problems 1 and 2 is a hierarchical two-step algorithm. The first step differs for the two problems and has as objective the generation of a reference path. The second step uses that reference path and leverages linear programming and a transformation of vehicle dynamics from the time into the spatial domain.The general approach to address both Problems 1 and 2 is a hierarchical two-step algorithm. The first step differs for the two problems and has as objective the generation of a reference path. The second step uses that reference path and leverages linear programming and a transformation of vehicle dynamics from the time into the spatial domain.
@@
Fig. 3: The kinematic bicycle model including the representation of the path-aligned coordinate system, see Rajamani (2011); Plessen et al. (2017a). The centre-of-gravity (CoG) of the vehicle is assumed to be located at the rear-axle and at a distance of $l>0$ to the steerable front-axle. Control inputs are velocity $v>0$ and steering angle $\delta\in[-\delta_{\text{max}},\delta_{\text{max}}]$. In the global position coordinate system the CoG-position is denoted by $(x,y)\in\mathbb{R}^{2}$ and the vehicle heading is $\psi\in[0,2\pi]$. This vehicle state can likewise be expressed in a path-aligned coordinate system, where $e_{\psi}\in[-\pi,\pi]$ and $e_{y}\in\mathbb{R}$ represent relative heading and lateral deviations with respect to references along the path centreline coordinate $s>0$. The radius of curvature at that coordinate is denoted by $\rho_{s}(s)>0$.
3 Theory
@@ 3.1 Spatial Domain Transformation
To illustrate the spatial domain transformation and for simplicity, the method is illustrated for a kinematic bicycle model in Fig. 3. Readers are referred to Plessen et al. (2017a) for the handling of more complex vehicle models.
This section summarises the bicycle model using the notation of Plessen et al. (2017a). Consider a global and a path-aligned coordinate frame within the $(x,y)$- and $(s,e_{y})$-plane, respectively, with path centreline coordinate $s\geq 0$ and lateral deviation $e_{y}\in\mathbb{R}$. Coordinates can be transformed by projecting point masses (vehicle’s centre of gravity) to piecewise-affine path segments. As a detail, it can be differentiated between the path centreline coordinate $s\geq 0$, and actual path length coordinate $\eta\geq 0$, indicating the distance traveled by the vehicle. Note that $\eta\neq s$, unless the vehicle is traveling perfectly along the path centreline. A trajetory can then be defined as $z(s)=\begin{bmatrix}e_{\psi}(s),~{}e_{y}(s)\end{bmatrix}$, with heading deviation $e_{\psi}(s)\in[0,2\pi]$ at coordinate $s$. The equivalent trajectory in the $(x,y)$-plane is defined by $\mathcal{X}(s)=\begin{bmatrix}x(s),~{}y(s),~{}\psi(s)\end{bmatrix}$. The classic nonlinear kinematic bicycle model (Rajamani (2011)) is
@@
assuming the centre of gravity to be located at the rear axle and $l$ denoting the wheelbase. The front-axle steering angle is $\delta$ and vehicle velocity is $v$. Let time and spatial derivatives be denoted by $\dot{x}=\frac{dx}{dt}$ and $x^{\prime}=\frac{dx}{ds}$, respectively. The spatial equivalent of Eqn. (1) is derived as follows. Assuming $\dot{e}_{\psi}=\dot{\psi}-\dot{\psi}_{s}$ where $\psi_{s}\in[0,2\pi]$ the path heading, $\dot{e}_{y}=v\sin(e_{\psi})$, $\dot{s}=\frac{\rho_{s}v\cos(e_{\psi})}{\rho_{s}-e_{y}}$ where $\rho_{s}$ denotes the radius of curvature, $e_{\psi}^{\prime}=\frac{\dot{e}_{\psi}}{\dot{s}}$ and $e_{y}^{\prime}=\frac{\dot{e}_{y}}{\dot{s}}$, then
@@
Importantly, Eqn. (2) is entirely independent of vehicle speed $v$. This is characteristic for kinematic models, but not the case for dynamic vehicle models (Plessen et al. (2017a)). Furthermore, note that Eqn. (2) has one state dimension less than Eqn. (1). To summarise, we define spatially-dependent state and control vectors as $z=\begin{bmatrix}e_{\psi},~{}e_{y}\end{bmatrix}$ and $u=\begin{bmatrix}\delta\end{bmatrix}$, respectively. Equation (2) is abbreviated by $z^{\prime}=f(z,u)$. Let a discretisation grid along the path centreline be defined by $\{s_{j}\}_{j=0}^{N}=\{s_{0},s_{1},\dots,s_{N}\}$. For the remainder of this paper, index $j$ refers to the spatial discretisation grid. The discretisation grid can be uniformly spaced. Additional grid points can be added. For a given set of references $\{e_{\psi,j}^{\text{ref}}\}_{j=0}^{N}$, $\{e_{y,j}^{\text{ref}}\}_{j=0}^{N}$ and $\{u_{j}^{\text{ref}}\}_{j=0}^{N-1}$ and an initial state, the linearised and discretised (zero-order hold) state dynamics of Eqn. (2) can be computed for $\{e_{\psi,j}\}_{j=0}^{N}$ and $\{e_{y,j}\}_{j=0}^{N}$ as a function of control inputs $\{\delta_{j}\}_{j=0}^{N-1}$.
@@
whereby the bounds, $\dot{\delta}^{\text{min}}$ and $\dot{\delta}^{\text{max}}$, in general, are nonlinear functions of the vehicle’s operating point and time-varying parameters. By applying the spatial coordinate transformation $\delta^{\prime}=\frac{\dot{\delta}}{\dot{s}}$, a discretisation, a change of variables, and assuming the bounds to remain constant over the planning horizon, then
@@
where $D_{s,j}=s_{j+1}-s_{j}$. Thus, the linear rate constraints in Eqn. (3) are rendered not only nonlinear, but also state-dependent and also velocity-dependent. This has two implications. First, to formulate linearly constrained optimisation problems, in general, the linearisation of Eqn. (4) is required. Depending on the quality of underlying reference trajectories, this may incur significant distortions. Second, while the discrete form of Eqn. (3) can always be guaranteed to be feasible (assuming a feasible initialisation), for Eqn. (4) this is not the case anymore. Consider two degrees of simplification of Eqn. (4). First, assume that the given reference trajectories permit small errors $e_{\psi,j}$ and $e_{y,j}$, and consequently approximate $\frac{(\rho_{s,j}-e_{y,j})}{\rho_{s,j}\cos(e_{\psi,j})}\approx 1$, thereby rendering Eqn. (4) state-independent, but maintaining velocity-dependent bounds. Thus, steering rate constraints still depend on $v_{j}$. Second, this velocity-dependency can be eliminated such that
@@
where, $v_{\text{ref}}>0$ is introduced as a reference velocity parameter choice. Note that the approximation in Eqn. (5) is the exact equivalent of Eqn. (4) for the zero-reference error setting $e_{y,j}=0$, $e_{\psi,j}=0$, $v_{j}=v_{\text{ref}}$. The formulation in Eqn. (5) bears the advantage of eliminating state dependency and is therefore the preferred form for spatial rate constraints. In practice, bounds are employed that are constant over the spatial planning horizon, but in general time-varying in any closed-loop receding horizon control-setting and dependent on the vehicle’s operating point.
To summarise this section, the transformation of time-dependent control rate constraints in Eqn. (3) to the path-aligned coordinate frame is not trivial and to be considered as the main disadvantage of a spatial-based system representation. The solution proposed here has opted for the simple state-independent form in Eqn. (5) for linearly constrained optimisation problems, and discussed the role of $v_{\text{ref}}>0$ as a transformation parameter.
Finally, an important analogy needs to be made between clothoid and continuous curvature path planning methods to further justify the above approach for the handling of rate constraints in the spatial domain. Continuous curvature and clothoid based path planning methods, such as in Sabelhaus et al. (2013), typically assume bounds on ’curvature’ and ’sharpness’ (the change in curvature) resulting in piecewise-affine curvature profiles along the planned path. Note that this implicitly also assumes an approximation of steering rate constraints.